# IM360: Textured Mesh Reconstruction for Large-scale Indoor Mapping with 360° Cameras

### v2

@@ Abstract
We present a novel 3D reconstruction pipeline for 360° cameras for 3D mapping and rendering of indoor environments. Traditional Structure-from-Motion (SfM) methods may not work well in large-scale indoor scenes due to the prevalence of textureless and repetitive regions. To overcome these challenges, our approach (IM360) leverages the wide field of view of omnidirectional images and integrates the spherical camera model into every core component of the SfM pipeline. In order to develop a comprehensive 3D reconstruction solution, we integrate a neural implicit surface reconstruction technique to generate high-quality surfaces from sparse input data. Additionally, we utilize a mesh-based neural rendering approach to refine texture maps and accurately capture view-dependent properties by combining diffuse and specular components. We evaluate our pipeline on large-scale indoor scenes from the Matterport3D and Stanford2D3D datasets. In practice, IM360 demonstrate superior performance in terms of textured mesh reconstruction over SOTA. We observe accuracy improvements in terms of camera localization and registration as well as rendering high frequency details. Project page: anonymousim360.github.io/
@@ 1 Introduction
Figure 1: This example compares two benchmark datasets: one captured with a conventional camera [11] and the other with an omnidirectional camera [6]. The blue wireframes indicate the camera locations. While the wide field of view of the omnidirectional camera significantly reduces the data acquisition time, it introduces challenges such as sparse views and occlusions. Our approach provides an effective solution to handle issues in sparsely scanned indoor environments and enables efficient reconstruction and rendering using omnidirectional cameras.
@@
  • We integrate classical texture mapping with neural texture fine-tuning through differentiable rendering, demonstrating improved rendering quality in sparsely scanned, in-the-wild indoor environments.
@@ 3 Our Method
Given a set of equirectangular projection (ERP) images, our goal is to reconstruct the geometry and texture of real-world scenes. In Fig. 2, our pipeline is composed of three primary components: Spherical Structure from Motion (SfM), geometry reconstruction, and texture optimization. To address challenges associated with limited overlap and frequent occlusions in indoor environments, our method leverages the wide field of view inherent in spherical imagery. We perform SfM utilizing a spherical camera model and dense features to estimate camera poses from sparsely scanned images in textureless regions. For 3D geometric reconstruction accurately, we converted ERP images into cubemaps and conducted training on the resulting perspective images. Then, we employ a neural model [62] that represents a signed distance function (SDF). After training the model, we extract the mesh using the Marching Cubes algorithm [38] and apply classical texture mapping [59] to initialize texture. Finally, we optimize the texture for diffuse, specular and small MLP components using differentiable rendering [34].
Figure 2: Overview of Our Pipeline. It is comprised of three steps: Spherical Structure from Motion (Sec. 3.1), Geometry Reconstruction (Sec. 3.2), and Texture Optimization (Sec. 3.3). We introduce a spherical SfM framework for estimating camera poses from omnidirectional images in challenging indoor environments. This is followed by geometry reconstruction to generate a 3D mesh, combined with texture mapping and neural texture optimization to enable realistic rendering of the entire scene.
Figure 3: Our Approach: In Section 3.1, spherical SfM is utilized to estimate the camera poses for omnidirectional images. Then, in Section 3.2, our method apply cubemap projection and estimate monocular depth and normal to train geometry $f_{sdf}$ representation using volumetric rendering. In Section 3.3, we initlize texture using classical texture mapping and jointly optimize diffuse, specular texture along with small MLP for modeling appearance.
@@ 3.1 Dense Matching and Structure from Motion
Large-scale indoor scenes often present significant challenges for camera localization when using conventional SfM methods such as COLMAP [50] or OpenMVG [44], due to numerous occlusions and textureless regions. To effectively reconstruct sparsely captured indoor scenes, we integrate a spherical camera model into every step of SfM pipeline, pipelining spherical dense matching, two-view geometry estimation, and optimization methods in a unified and complete 3D reconstruction framework. Instead of using the traditional pinhole camera model and normalized coordinates, we build our pipeline based on the unit bearing vector representation $u\in\mathcal{S}^{2}$ and a spherical projection mapping $S$, where $S$ projects as $S:P\rightarrow u,\quad\mathbb{R}^{3}\mapsto\mathbb{S}^{2}$. We provide a description of our SfM core components with unified spherical representations.
Spherical Dense Matching In our approach, we borrow EDM [1], which extends Dense Kernelized Matching (DKM) [14] with a spherical camera model. DKM formulates the feature matching problem as a form of Gaussian Process Regression (GPR) problem, assuming a mapping $f:\varphi\rightarrow\mathcal{X}$ as a Gaussian process to be estimated, where $\varphi$ and $\mathcal{X}$ denotes feature descriptors and pixel coordinate embeddings, respectively. Given paired images $A$ and $B$, EDM predicts the posterior Gaussian distribution of spherical coordinate embeddings $\mathcal{U}_{A}\,|\,\varphi_{A},\mathcal{U}_{B},\varphi_{B}\sim\mathcal{N}(\mu,\Sigma)$ conditioned on the observations $(\mathcal{U}_{B},\varphi_{B})$ from the image $B$, which is formulated as,
$\mu$ $=K_{AB}(K_{BB}+\sigma_{n}^{2}I)^{-1}\mathcal{U}_{B},$ (1)
$\Sigma$ $=K_{AA}-K_{AB}(K_{BB}+\sigma_{n}^{2}I)^{-1}K_{BA},$ (2)
where $\mu$ is refined through CNN-based warp refiners [1, 14].
Spherical Two-view Geometry Estimation It can be easily proven that the traditional two-view epipolar constraint for normalized coordinates can be extended to unit bearing vectors on spherical images,
$u_{1}^{T}Eu_{2}=0\;\text{with}\;E=R[t]_{\times}.$ (3)
@@
Spherical Bundle Adjustment By representing each feature track observation as a bearing vector, conventional Bundle Adjustment (BA) can be reformulated in a manner similar to the perspective case, as follows:
$L=\sum_{i}\sum_{j}\rho(||S(P_{j};R_{i},t_{i})-u_{ij}||^{2}),$ (4)
[…]
@@ 3.2 Geometric Reconstruction
To address visual localization and mapping in textureless and highly occluded sparsely scanned indoor scenes, we use neural rendering techniques [64, 62] for surface reconstruction instead of traditional 3D reconstruction methods such as multi-view stereo (MVS) [51]. Many methods leverage monocular geometric priors [64, 62] and demonstrate higher robustness in textureless and sparsely covered areas. However, in terms of ERP images there are no robust methods for zero-shot monocular depth and normal estimation. We also obsered that OmniSDF [32], which learns signed distance functions (SDF) from equirectangular projection (ERP) images, failed to converge when applied to large-scale indoor datasets.
To overcome these limitations, we transformed ERP images into cubemaps, estimated depth $D$ and normal maps $N$ from all six perspective images using a pretrained Omnidata model [16], and utilized these geometric cues to train neural surface reconstruction effectively. Following the DebSDF [62], we jointly train two MLPs using the differentiable volumetric rendering, (i) $f_{sdf}$, which represents the scene geometry as a signed distance function, and (ii) $f_{color}$, a color network. The training process [62] incorporates a combination of losses, including color reconstruction loss $L_{\text{rgb}}=\sum_{r\in R}||\hat{C}(r)-C(r)||_{1}$, Eikonal loss [21] $L_{\text{eikonal}}=\sum_{x\in\chi}(||\triangledown f_{sdf}(x)||_{2}-1)^{2}$ , and depth and normal losses, The latter are derived from prior geometric cues by comparing the rendered depth $\hat{D}(t)$ and normals $\hat{N}(t)$ with the corresponding prior depth $D$ and normals $N$. Color image $\hat{C}$ is volumetrically rendered by ray marching $\hat{C}=\sum_{i\in I}\alpha_{i}\,C_{i}\,T_{i}$ along with $\hat{D}$ and $\hat{N}$. We then utilize the learned SDF ($f_{sdf}$ evaluated over a uniform grid) to extract a mesh $M$ using the Marching Cubes algorithm [38].
@@ 3.3 Texture Optimization
Following TMO [9], we parameterize the texture map and leverage a differentiable rasterizer $\mathcal{R}$ [34] to render an image from a given viewpoint $P$. Unlike TMO [9], which is limited to representing diffuse textures and cannot capture view-dependent effects, our approach overcomes this limitation. The diffuse texture $K_{d}\in\mathbb{R}^{3}$ is directly converted into an RGB image. Additionally, we initialize a specular feature $K_{s}\in\mathbb{R}^{3}$ and employ a small MLP $f_{s}$ as a fragment shader. This MLP takes the specular feature $K_{s}$ and viewing direction $v$ as inputs to compute the specular color as follows:
@@ 4.1 Experiment Settings
The Matterport3D dataset [6] consists of 90 indoor scenes encompassing a total of 10,800 panoramic images. Following the EDM approach [1], we define the ground truth camera poses of these images. From the validation and test sets of the official benchmark split, we selected three scenes each, ensuring that these scenes are entirely indoors or composed of levels no higher than the second floor. Since dense matching requires overlapped image pairs as input, we assume that during 360-degree camera scanning, the information about which room each camera belongs to is known. To utilize this information effectively, we leveraged the $house\_segmentation$ labels available in the Matterport3D dataset. Images taken within the same room were selected as matching pairs, while for transitions between different rooms, only one image pair with the highest overlap was linked. For EDM [1], the ERP images were resized to $640\times 320$, whereas images for other feature matching methods were resized to $1024\times 512$.
@@ 4.2 Experimental Results
4.3 Ablation Study
We observed that when initial poses are determined through two-view geometry without utilizing the method proposed by Solarte et al. [52], the registration performance significantly deteriorates, resulting in fragmented SfM models. Furthermore, as shown in Table 1 and 2, by modifying the feature matching in COLMAP-based methods and comparing the registration and pose accuracy, we confirmed that our approach demonstrates superior robustness. In table 3, IM360* only finetunes diffuse texture without utilizing specular features. Without texture optimization stage outlined in Sec 3.3, our rendering quality is equivalent to that of TexRecon [59]. Although we apply texture optimization soley to the diffuse texture map, our method achieves a substantial improvement in rendering quality, with a 2.7 PSNR increase. Additionally, our final approach, which combines both diffuse and specular colors, provides a further 0.8 PSNR gain by accounting for view-dependent components.
@@ 5 Conclusion, Limitations and Future Work
We introduce a novel framework for textured mesh reconstruction for omnidirectional cameras in sparsely scanned indoor environments. To address the challenges associated with textureless regions and sparse viewpoints, we propose a spherical Structure-from-Motion (SfM) that integrates spherical camera models into all the core steps of SfM. Additionally, we reconstruct texture maps and leverage differentiable rendering to enhance its quality, improving the rendering performance in sparse-view scenarios. However, our approach has certain limitations. First, the performance of dense matching depends on correct matching pairs. Therefore, if we lack an appropriate image retrieval technique, we must rely on covisibility annotations during the scanning process. Furthermore, our texture mapping method is based on a mesh rasterization approach, making the rendering quality highly sensitive to the accuracy of the underlying mesh geometry. Consequently, any inaccuracies in the mesh can result in rendering artifacts. In the future, we may explore ERP image retrieval methods to extract covisible image pairs and the joint optimization of geometry and texture to refine geometric artifacts.

### v3

# IM360: Large-scale Indoor Mapping with 360 Cameras

@@ Abstract
We present a novel 3D mapping pipeline for large-scale indoor environments. To address the significant challenges in large-scale indoor scenes, such as prevalent occlusions and textureless regions, we propose IM360, a novel approach that leverages the wide field of view of omnidirectional images and integrates the spherical camera model into the Structure-from-Motion (SfM) pipeline. Our SfM utilizes dense matching features specifically designed for 360 images, demonstrating superior capability in image registration. Furthermore, with the aid of mesh-based neural rendering techniques, we introduce a texture optimization method that refines texture maps and accurately captures view-dependent properties by combining diffuse and specular components. We evaluate our pipeline on large-scale indoor scenes, demonstrating its effectiveness in real-world scenarios. In practice, IM360 demonstrates superior performance, achieving a 3.5 PSNR increase in textured mesh reconstruction. We attain state-of-the-art performance in terms of camera localization and registration on Matterport3D and Stanford2D3D. Project page: https://jdk9405.github.io/IM360/
@@ 1 Introduction
Dataset # Scenes # Images Floor Space (m2) View Density (images/m2)
Stanford2D3D [1] 6 1,413 (Omnidirectional) 6020 0.23
Matterport3D [5] 90 10,800 (Omnidirectional) 46,561 0.23
ScanNet [12] 707 2,492,518 (Conventional) 34,453 72.35
Figure 1: While the wide field of view of the omnidirectional camera significantly reduces the data acquisition time, it introduces challenges such as sparse views and occlusions. According to the Table, Matterport3D [5] and Stanford2D3D [1] are captured with significantly lower view density compared to ScanNet [12]. Our approach provides an effective solution to handle issues in sparsely scanned indoor environments and enables efficient reconstruction and rendering using omnidirectional cameras.
@@
  • We integrate classical texture mapping with neural texture fine-tuning through differentiable rendering, demonstrating improved rendering quality, in sparsely scanned, in-the-wild indoor environments.
@@ 3 Our Method
Given a set of equirectangular projection (ERP) images, our goal is to reconstruct the geometry and texture of real-world scenes. In Fig. 2, our pipeline is composed of three primary components: Spherical Structure from Motion (SfM), geometry reconstruction, and texture optimization. To address challenges associated with limited overlap and frequent occlusions in indoor environments, our method leverages the wide field of view inherent in spherical imagery. We perform SfM utilizing a spherical camera model and dense features to estimate camera poses from sparsely scanned images in textureless regions. However, since a naive combination of SfM and dense feature matching is not directly feasible, we refine this approach to ensure compatibility. For 3D geometric reconstruction accurately, we converted ERP images into cubemaps and conducted training on the resulting perspective images. Then, we employ a neural model [74] that represents a signed distance function (SDF). After training the model, we extract the mesh using the Marching Cubes algorithm [43] and apply classical texture mapping [70] to initialize texture. Finally, we optimize the texture for diffuse, specular and small MLP components using differentiable rendering [38].
Figure 2: Overview of Our Pipeline. It is comprised of three steps: Spherical Structure from Motion (Sec. 3.1), Geometry Reconstruction (Sec. 3.2), and Texture Optimization (Sec. 3.3). In Section 3.1, spherical SfM is utilized to estimate the camera poses for omnidirectional images in challenging indoor environments. Then, in Section 3.2, our method apply cubemap projection and estimate monocular depth and normal to train geometry $f_{sdf}$ representation using volumetric rendering. In Section 3.3, we initlize texture using classical texture mapping and jointly optimize diffuse, specular texture along with small MLP for modeling appearance.
@@ 3.1 Dense Matching and Structure from Motion
Large-scale indoor scenes often pose significant challenges for camera localization using conventional SfM methods such as COLMAP [57] or OpenMVG [49], due to numerous occlusions and textureless regions. To effectively reconstruct sparsely captured indoor scenes, we integrate a spherical camera model into every step of SfM pipeline, combining spherical dense matching, two-view geometry estimation, and optimization methods into a unified, comprehensive 3D reconstruction framework. We construct our pipeline based on the unit bearing vector $u\in\mathbb{S}^{2}$,
$u=\left(\sin(\theta)\cos(\phi),\ \sin(\phi),\ \cos(\theta)\cos(\phi)\right)$ (1)
where $\theta\in[-\pi,\pi]$ and $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$.
Spherical Dense Matching Performing spherical SfM requires feature matching as a preliminary step. Several feature matching methods specifically designed for 360images have been proposed, such as SPHORB [78], SphereGlue [20], and EDM [34]. Alternatively, feature matching can also be performed independently on each face of a cubemap generated from a 360image. Among these approaches, dense feature matching [34] recently demonstrated superior feature matching performance for 360° images, especially in sparse scenes with large viewpoint changes. However, since dense matching methods do not explicitly provide feature descriptors, constructing feature tracks of multiple views for SfM becomes challenging. Inspired by [54, 26], we apply the additional strategies for match quantization by rounding [6] and merging [59] techniques, to ensure multiview match consistency.
Spherical Two-view Geometry Estimation It can be formally proven that the traditional two-view epipolar constratint for normalized coordinates extends to unit bearing vectors in spherical images [56],
$u_{1}^{T}Eu_{2}=0\;\text{with}\;E=R[t]_{\times},$ (2)
where $E\in\mathbb{R}^{3\times 3}$ is the essential matrix, which can be decomposed into the camera rotation $R\in\text{SO(3)}$ and the translation $t\in\mathbb{R}^{3}$. Solarte et al. [61] employed the 8-point algorithm to solve spherical two-view geometry using the Direct Linear Transform (DLT) framework [25]. We adopt this two-view geometry estimation method with additional normalization [61] to enhance numerical stability during Singular Value Decomposition (SVD) computation. Using those initial pose with rotation $R$ and translation $t$, we activate incremental triangulation and bundle adjustment under the sphere camera model.
Spherical Bundle Adjustment By representing each feature track observation as a bearing vector $u$, conventional Bundle Adjustment (BA) can be reformulated in a manner similar to the perspective case, as follows:
@@
where $\rho$ denotes the robust loss with a soft L1 function in local bundle adjustment, while global bundle adjustment is performed without this robustifier. $\Pi(\cdot)$ represents the spherical projection $\Pi:P\rightarrow u$, $\mathbb{R}^{3}\mapsto\mathbb{S}^{2}$, which maps the 3D points $X_{j}$ from triangulation to unit bearing vectors, given the camera poses $R_{j}\in\text{SO(3)}$ and $t_{j}\in\mathbb{R}^{3}$.
@@ 3.2 Geometric Reconstruction
Figure 3: Our method, which applies cubemap projection and trains neural surface reconstruction using multiple perspective images, achieves highly accurate geometry reconstruction. In contrast, neural surface reconstruction designed for ERP images (OmniSDF [36]) and Gaussian Splatting-based surface reconstruction (SuGaR [24] and VCR-GauS [7]) fail to generate a clean mesh.
To address visual localization and mapping in textureless and highly occluded sparsely scanned indoor scenes, we use neural rendering techniques [76, 74] for surface reconstruction instead of traditional 3D reconstruction methods such as multi-view stereo (MVS) [58]. Many methods leverage monocular geometric priors [76, 74] and demonstrate higher robustness in textureless and sparsely covered areas. However, in terms of ERP images there are no robust methods for zero-shot monocular depth and normal estimation. As shown in Fig. 3, we also observed that OmniSDF [36], which learns signed distance functions (SDF) from equirectangular projection (ERP) images, failed to converge when applied to large-scale indoor datasets. Despite incorporating geometric priors [17], Gaussian Splatting-based surface reconstruction [24, 7] still fails to produce reasonable geometry, resulting in a polygonal soup.
To address these limitations, we first project ERP images onto a cubemap representation and then apply a volumetric rendering approach to estimate the signed distance function (SDF). Since most existing mesh rendering techniques—except for volumetric methods leveraging monocular geometric priors [76, 74]—struggle in sparse indoor scenes, we employ DebSDF [74] for geometric mesh reconstruction and integrate the Marching Cubes algorithm [43] to generate a high-fidelity mesh.
@@ 3.3 Texture Optimization
Inspired by TMO [10], we parameterize the texture map and leverage a differentiable rasterizer $\mathcal{R}$ [38] to render an image from a given viewpoint $P$. While TMO [10] is constrained to representing only diffuse textures and fails to capture view-dependent effects, our approach surpasses these limitations by enabling a more comprehensive and physically accurate representation. The diffuse texture $K_{d}\in\mathbb{R}^{3}$ is directly converted into an RGB image. Additionally, we initialize a specular feature $K_{s}\in\mathbb{R}^{3}$ and employ a small MLP $f_{s}$ as a fragment shader. This MLP takes the specular feature $K_{s}$ and viewing direction $v$ as inputs to compute the specular color as follows:
@@ 4.1 Experiment Settings
The Matterport3D dataset [5] consists of 90 indoor scenes encompassing a total of 10,800 panoramic images. Following the EDM approach [34], we define the ground truth camera poses of these images. From the validation and test sets of the official benchmark split, we selected three scenes each, ensuring that these scenes are entirely indoors or composed of levels no higher than the second floor. Since dense matching requires overlapped image pairs as input, we assume that during 360-degree camera scanning, the information about which room each camera belongs to is known. To utilize this information effectively, we leveraged the house_segmentationhouse\_segmentation labels available in the Matterport3D dataset. Images taken within the same room were selected as matching pairs, while for transitions between different rooms, only one image pair with the highest overlap was linked. For EDM [34], the ERP images were resized to $640\times 320$, whereas images for other feature matching methods were resized to $1024\times 512$.
@@ 4.2 Experimental Results
Figure 7: Localization results in multi-floor scenes of Matterport3D. Multi-floor scenes pose significant challenges for visual localization due to narrow corridors and staircases connecting different floors. Despite these difficulties, IM360 demonstrates the ability to estimate poses, ensuring robust localization in large-scale indoor environments.
4.3 Additional Experiments and Ablation Study
We observed that when initial poses are determined using two-view geometry without applying the method proposed by Solarte et al. [61], registration performance significantly deteriorates, resulting in fragmented SfM results. Furthermore, as shown in Table 1 and 2, by modifying the feature matching in COLMAP-based methods and comparing the registration and pose accuracy, we confirmed that our approach demonstrates superior robustness. We also evaluate our SfM method on multi-floor scenes, as illustrated in Fig. 7. Our spherical SfM exhibits robust performance in multi-floor scenes, highlighting the advantages of leveraging spherical cameras within our proposed pipeline for large-scale indoor environments.
In table 3, IM360* only finetunes diffuse texture without utilizing specular features. Without texture optimization stage outlined in Sec 3.3, our rendering quality is equivalent to that of TexRecon [70]. Although we apply texture optimization soley to the diffuse texture map, our method achieves a substantial improvement in rendering quality, with a 2.7 PSNR increase. Additionally, our final approach, which combines both diffuse and specular colors, provides a further 0.8 PSNR gain by accounting for view-dependent components.
@@ 5 Conclusion, Limitations and Future Work
We introduce a novel framework for textured mesh reconstruction for omnidirectional cameras in sparsely scanned large-scale indoor environments. To tackle the challenges posed by textureless regions and sparse viewpoints, we propose a spherical Structure-from-Motion (SfM) that incorporates dense feature matching and integrates spherical camera models. Additionally, we reconstruct texture maps and leverage differentiable rendering to enhance its quality, improving the rendering performance in sparse-view scenarios. One limitation of our approach is its reliance on human annotation for determining matching pairs. While 360 image scanning scenarios typically involve a small number of captures, making manual annotation feasible, integrating an automated image retrieval system would be a more scalable solution. For geometric reconstruction, we employ cubemap projection to transform 360 images into perspective views and utilize state-of-the-art volumetric rendering methods. However, this approach has limited capacity for reconstruction, as the quality deteriorates when the scene exceeds a certain scale. In future work, we aim to further refine geometric reconstruction and enhance the overall pipeline for large-scale 360 mapping.
[…]
Supplementary Material
In this supplementary material, we provide additional experimental results, highlighting the superior performance of our method compared to existing approaches. Appendix 6 shows the qualitative results of our custom dataset to demonstrate the practicality of our method. Table 4 present the sensor configurations of publicly available datasets containing indoor scenes. Since Matterport3D [5] and Stanford2D3D [1] offer sparsely scanned panoramic images in large-scale environments, our experiments primarilty focus on these two datasets. Appendix 8.1 highlights the advantages of the spherical camera model and the dense matching algorithm for indoor reconstruction. Appendix 8.2 describes the details and justifies the model choices for geometry mesh reconstruction. Appendix 8.3 demonstrates the effectiveness of our novel texturing method.
Figure 9: Qualitative results of the custom dataset. We captured 20 images of a house using the commerical 360 sensor, Insta360.
Figure 10: Overview of Neural Surface Reconstruction. In this work, we follow DebSDF [74], which estimates an implicit surface representation by utilizing volumetric rendering and monocular geometric priors. This approach refines the underlying structure and enhances surface details, leading to more accurate 3D reconstructions compared to other recent methods.
Dataset Environment Floor Space (m2) # Perspective Images # Panoramic Images Annotation
Stanford2D3D [1] 6 large indoor areas 6020 70,496 1,413 Laser Scanner
Matterport3D [5] 2056 rooms (90 scenes) 46,561 194,400 10,800 Laser Scanner
ScanNet [12] 707 small rooms 34,453 2,492,518 - RGB-D
7 Scenes [60] 7 small rooms - 43,000 - RGB-D
TUM Indoor [29] building (7 floors) 16,341 48,974 - Laser Scanner
TUM-LSI [71] building (5 floors) 5,575 1,095 - Laser Scanner
InLoc [65] building (5 floors) 10,370 10328 - Laser Scanner / RGB-D
Baidu [63] mall 9,179 2,078 - Laser Scanner
NAVER LABS [39] mall and metro 53,036 136,783 - Laser Scanner
Table 4: Among large-scale indoor datasets, Matterport3D [5] and Stanford2D3D [1] provide 360 panoramic images, unlike other datasets that primarily rely on perspective images. Compared to perspective images, 360 images significantly reduce the number of captures required to cover a scene. However, this sparse and reduced number of images introduces various challenges for visual localization and mapping pipelines. To address these challenges, we integrate spherical SfM, geometric reconstruction, and texture optimization techniques into our approach.
6 Real World Application
To demonstrate the versatility of our proposed spherical Structure from Motion, we tested it on a custom dataset. We directly captured 20 omnidirectional images in a 56 m2 house using an Insta360 camera. Figure 9 presents all registered cameras obtained through spherical SfM along with the 3D reconstruction results.
7 Technical Details
7.1 Spherical Structure from Motion
Spherical Dense Matching: We provide details on spherical dense matching. Dense matchers estabilsh pixel-wise correspondences and sample reliable matches using confidence scores. Ideally, if the network is well-trained, unreliable matches are filtered by confidence, and further refined through geometric filtering during SfM. Furthermore, following the detector-free setting [59, 26], we quantize each 2D match location onto a grid to enhance the consistency of closely spaced subpixel matches. Specifically, each coordinate $x$ is rounded to the nearest grid point using the formula, $\left\lfloor\frac{x}{r}\right\rceil\cdot r$ where $r$ denotes the grid cell size and $\left\lfloor\cdot\right\rceil$ represents the rounding operator. This quantization step combines nearby matches into the same grid cell, effectively merging redundant or noisy matches into a single representative location.
Spherical Two-view Geometry Estimation: We provide additional details on spherical two-view geometry estimation in the manuscript. Following Solarte et al. [61], we adopt the eight-point algorithm (8-PA) to estimate the essential matrix from corresponding points. Unlike the conventional 8-PA [25], which uses normalized image coordinates, we replace them with unit bearing vectors on a unit sphere via spherical projection. Thus, a minimum of $n\geq 8$ bearing vector correspondences is required to estimate the essential matrix. We reformulate Equation (2) from the manuscript as a least-squares problem:
$A[E]_{v}=0$ (6)
Here, $A$ is an $n\times 9$ matrix formed by stacking the Kronecker products of corresponding bearing vectors as $A_{i}=u^{i}_{1}\otimes u^{i}_{2}$, and $E]_{v}$ is a vector obtained by row-wise concatenation of the essential matrix entries. The solution to Equation 6 is obtained using the Direct Linear Transformation (DLT) method [25], from which the essential matrix can be recovered.
Image Pair Selection: We do not consider image pair selection a core component of our pipeline and do not incorporate image retrieval methods such as NetVLAD, as we regard this as a separate research topic. We intentionally avoid using automatic annotation to ensure accurate evaluation of other components (Sec3.1), especially since human annotation is relatively easy in sparse view settings.
7.2 Geometric Reconstruction
[…]
The depth and normal losses are derived from prior geometric cues by comparing the rendered depth $\hat{D}(t)$ and normals $\hat{N}(t)$ with the corresponding prior depth $D$ and normals $N$ from Omnidata [17]. Color image $\hat{C}$ is volumetrically rendered by ray marching $\hat{C}=\sum_{i\in I}\alpha_{i}\,C_{i}\,T_{i}$ along with $\hat{D}$ and $\hat{N}$. We then utilize the learned SDF ($f_{sdf}$ evaluated over a uniform grid) to extract a mesh $M$ using the Marching Cubes algorithm [43].
Figure 11: Examples of VCR-Gauss [7] and DebSDF [74]
8 More Experimental Results
8.1 Spherical Structure from Motion
Due to page limitations, we present our spherical structure from motion results in the supplementary material: 1) OpenMVG: An open-source SfM pipeline that supports spherical camera models [49]. 2) SPSG COLMAP: SuperPoint [13] and SuperGlue [55] are used with cubemap and equirectangular projection. 3) DKM COLMAP: This method leverages DKM [15] to establish dense correspondences, utilizing cubemap and equirectangular projection. 4) SphereGlue COLMAP: SuperPoint [13] with a local planar approximation [14] and SphereGlue [20] are utilized to mitigate distortion in ERP images. The experimental results discussed in the main paper for Matterport3D [5] and Stanford2D3D [1] are shown in Fig. 12 and Fig. 13, respectively.
8.2 Geometric Reconstruction
Figure 11 presents the geometry reconstruction results of VCR-Gauss [7] for comparison. VCR-Gauss is a Gaussian Splatting-based surface reconstruction method that utilizes monocular geometry priors, similar to DebSDF [74]. However, in our experiment, we completely fail to train this model, resulting in a polygonal soup.
8.3 Texture Map Optimization
We compare our method with several recent rendering approaches, including TexRecon [70], SparseGS [75], and ZipNeRF [3]. Our method outperforms these approaches by delivering higher frequency details and producing seamless texture maps. The results of the textured mesh and rendering are shown in Fig. 14 and Fig. 15 - 18.
8.4 Acknowledgements
This work was supported in part by ARO Grants W911NF2310046, W911NF2310352, and U.S. Army Cooperative Agreement W911NF2120076.
Figure 12: Qualitative Comparison of SfM results on Matterport3D. While other approaches failed to achieve pose registration, our method successfully estimates poses by leveraging the spherical camera model and dense matching.
Figure 13: Qualitative Comparison of SfM results on Stanford2D3D. While other approaches failed to achieve pose registration, our method successfully estimates poses by leveraging the spherical camera model and dense matching.
Figure 14: Qualitative Comparisons of Textured Mesh Results on Matterport3D. A comparison between TexRecon [70] and ours shows that our method effectively reduces noise in the texture maps, leading to improved visual quality and detail.
Figure 15: Qualitative Comparisons with Existing Methods. Our method can render high frequency details and results in lower noise.
Figure 16: Qualitative Comparisons with Existing Methods. Our method can render high frequency details and results in lower noise.
Figure 17: Qualitative Comparisons with Existing Methods. Our method can render high frequency details and results in lower noise.
Figure 18: Qualitative Comparisons with Existing Methods. Our method can render high frequency details and results in lower noise.